OpenStudy (anonymous):

Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the team have?

OpenStudy (anonymous):

wins (w); ties (T) = W - 9 2W + (W -9) = 60 3W - 9 = 60 3W = 69 W = 23 Therefore.... 23 wins and 14 ties (23-9=14)

OpenStudy (anonymous):

thank you :D